A baseball team has home games on Thursday and Saturday. The two games together earn $4404.00 for the team. Thursday's game generates $376.00 less than Saturday's game. How much money was taken in at each game?
Answer by macston(5194)

.
T = amount on Thursday; S = amount on Saturday = T+$376
.
T+S=$4404
T+(T+$376)=$4404
2T=$4028
T=$2014
ANSWER 1: The Thursday game took in $2014.
.
S=T+$376
S=$2014+$376
S=$2390
ANSWER 2: The Saturday game took in $2390.
.
CHECK:
S+T=$4404
$2390+$2014=$4404
$4404=$4404
.